y varies directly with x if y=16 when x=48
A: find the constant of variation,k, and write an equation for the direct variation.
B: find x when y=-27

Answer by Mathtut(3670)

in direct proportions if y varies directly with x then 16/48=1/3
1. so the constant k is 1/3 and all y/x terms must be in that proportion
2.so 16/48=-27/x so 16x=-27(48)---->16x=-1296 so x=-81 you will see that if you take -27/-81 you will have a 1/3 ratio if you simplify